OpenStudy (anonymous):

A sound wave has a velocity of 350 meters/second and a frequency of 430 Hertz. What is it's wavelength? A. 0.01 B. 0.60 C. 0.81 D. 0.08

OpenStudy (anonymous):

Frequency= Speed / Wavelength Wavelength = Speed / Frequency = 350/430 =0.81 (C)

OpenStudy (anonymous):

@Kaylalove Is this the equation f=c/lambda????I think it applies to EM waves only

OpenStudy (s3a):

@Krishnadas, I'm pretty sure that it's f = v/lambda and you replace v with c when it's light. In other words, I believe that all waves use these equations.

OpenStudy (anonymous):

@s3a ....Hmm thats confusing!!! :D